OpenStudy (anonymous):

Solve x + 3y = 9 3x – 3y = –13

OpenStudy (ajprincess):

add both the eqns Then u get 4x=-4 divide by 4 x=-1 Substitute this in the first eqn -1+3y=9 solve for y.

OpenStudy (cwrw238):

add 2 equations to eliminate y solve for x then substitute this value for x in first equation to find y

OpenStudy (anonymous):

We can just add them together since the y-values will cancel out. x+3y=9 3x-3y=-13 -------------- 4x=-4 Now we divide by 4 and get x=-1. Now we put this into one of the equations, doesn't matter which one it is, and solve for the y-value. -1+3y=9 Now we add bye one to both sides. 3y=10 and now we divide by 3 y=3.33333...... So our answer is (-1,3.3...)

OpenStudy (kaiz122):